A charge of 5C charge is placed at origin and a charge of 3C find at 2meter distance find force due to charge on each other
1
Expert's answer
2020-09-22T17:08:46-0400

We can find the force between the two charges from the Coulomb's law:


"F = \dfrac{kq_1q_2}{r^2},"

here, "F" is the force between the two charges, "k = 9 \cdot 10^9 \ \dfrac{N \cdot m^2}{C^2}" is Coulomb's constant, "q_1 = 5 \ C" is the magnitude of the first charge, "q_2 = 3 \ C" is the magnitude of the second charge,

"r = 2 \ m" is the distance between the charges.

Then, we can calculate the force between the two charges:


"F = \dfrac{9 \cdot 10^9 \ \dfrac{N \cdot m^2}{C^2} \cdot 5 \ C \cdot 3 \ C}{(2 \ m)^2} = 33.75 \cdot 10^9 \ N."

Answer:

"F = 33.75 \cdot 10^9 \ N."